OpenStudy (anonymous):

Which fraction is greater than -7/15 ? A. -5/9 B. -6/11 C. -1/2 D. -4/9

OpenStudy (anonymous):

-7/15 equals -.49999999

OpenStudy (anonymous):

So which number is greater than that?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

You can solve this problem without using a calulator. -7/15 is greater than 50% That measn our answer MUST BE BELOW 50% because they are all negative.

OpenStudy (anonymous):

A is wrong because -5/9s is greater than 50%

OpenStudy (anonymous):

B is wrong because -6/11s is greater than 50%

OpenStudy (anonymous):

C is wrong because -1/2 is exactly 50%

OpenStudy (anonymous):

D

OpenStudy (anonymous):

We only have one answer left. Yes you are correct D is correct because is is less that 50% (-4/9)=-.44444

OpenStudy (charlotte123):

It is D :D just checked on my calculator and went step by step ^_^ -7/15 = -0.46.... when -4/9 = 0.44..... ^_^

OpenStudy (gamer56):

C. was right and thats exactly what i put :D